# Optimization via a Control-Centric Framework

### v2

@@ V-B Convex-Concave Case
By Sion’s minimax theorem [34], a saddle point $(\mathbf{x}^{\star},\mathbf{y}^{\star})$ exists. Along the dynamics (55), the derivative of $V$ satisfies
@@
Classical minimax dynamics have established asymptotic or exponential convergence under convex-concave assumptions [12, 13]. Our framework recovers these results as special cases and further guarantees FT, FxT, or PT convergence by an appropriate choice of the decay law.
@@ V-C Strongly Convex-Strongly Concave Case
Similarly to the constrained optimization case, when strong convexity/concavity holds we get automatic nonsingularity of Assumption III.2.
@@
Classical constrained minimax dynamics achieve exponential convergence for convex-concave problems [12, 13]. Recent FT/FxT solutions address unconstrained settings [20, 18, 33]. Theorem V.1 and Corollary V.2 extend FT/FxT/PT guarantees to constrained convex-concave minimax problems.
This section establishes convergence for constrained convex-concave minimax problems according to a chosen convergence rate, extending beyond existing asymptotic or exponential guarantees in the literature. These results illustrate how the OLF framework generalizes across problem classes with shared feasibility constraints. In the next section, we turn to the GNE seeking problems, where multiple agents interact strategically under such shared constraints.
@@ VI Generalized Nash Equilibrium Seeking Problems
Generalized Nash equilibrium (GNE) seeking problems arise when agents share coupling constraints, such as common resources or safety requirements, while pursuing individual objectives. These problems have been extensively studied from both the variational inequality and optimization perspectives; see the comprehensive survey in [35]. Classical reformulations based on Nikaido-Isoda functions and their regularized variants [36] provide optimization-based characterizations of (normalized) GNE. Algorithmic approaches have primarily relied on penalty and augmented Lagrangian methods, which embed the constraints into the agents’ costs. Representative contributions include penalty reformulations [37], augmented Lagrangian frameworks [38], and, more recently, continuous-time penalty dynamics that establish asymptotic or exponential convergence [39].
@@ VI-A Problem statement and stationarity vector
Consider $N$ players with decision vectors $\mathbf{x}_{i}\in\mathbb{R}^{n_{i}}$. A profile $\bar{\mathbf{x}}=\operatorname{col}(\bar{\mathbf{x}}_{1},\dots,\bar{\mathbf{x}}_{N})$ with $n=\sum_{i}n_{i}$ is a GNE if it is feasible for the shared constraints and no player can reduce its cost by a unilateral deviation that also respects those shared constraints; namely,
@@
with $\mathcal{X}_{i}(\mathbf{x}_{-i}):=\{\mathbf{x}_{i}\mid\mathbf{A}\mathbf{x}=\mathbf{b},\ \mathbf{g}(\mathbf{x})\leq\mathbf{0}\}$ where $\mathbf{A}\in\mathbb{R}^{q\times n}$, $\mathbf{b}\in\mathbb{R}^{q}$, and $\mathbf{g}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ is $C^{1}$. This differs from a standard Nash equilibrium because each player’s feasible set depends on the others via the shared constraints. In convex settings, equilibria for which all players share the same multipliers $(\boldsymbol{\lambda},\boldsymbol{\mu})$ for the joint constraints are called variational GNE (v-GNE); these coincide with solutions of a single KKT system [40] and are the natural target for our OLF design.
@@ VI-B Main result
Note that Assumption III.2 holds automatically in this setting.
@@
Existence and uniqueness of the variational equilibrium follow from strong monotonicity of $\mathcal{G}$ and convexity of the shared feasible set, as established by variational inequality theory [35, Sec. 3.2]. By construction of $\mathbf{u}(\mathbf{z},t)$, the OLF $V(\mathbf{z})$ satisfies the chosen decay law, hence $V(\mathbf{z}(t))\!\to\!0$ with the corresponding temporal guarantee, and thus $\mathbf{S}(\mathbf{z}(t))\!\to\!\mathbf{0}$.
@@ VI-C Numerical Example: Cournot GNE with Shared Constraints
We illustrate Thm. VI.2 with a Cournot competition game [41]. Consider $N=4$ firms and $M=2$ markets. Each firm selects a production vector $\mathbf{x}_{k}\in\mathbb{R}^{2}_{\geq 0}$, and the aggregate supply is $C\mathbf{x}$. Market prices follow a linear inverse demand
@@
It is also worth noting that the observation of Remark IV.7 carries over unchanged here: since the Lyapunov function is built directly from the stacked stationarity vector, driving $V(z)$ below a given tolerance automatically certifies that each optimality component (stationarity, feasibility, and complementarity) is satisfied to the same numerical accuracy. This single scalar measure thus provides a unified guarantee for convergence to the v-GNE.
@@ VII Discussion About Discrete-Time Implementation
This section discusses the discrete-time realization of our continuous-time designs. Several recent directions inform the discrete-time design of continuous-time optimization dynamics. Discrete-gradient (DG) methods construct a step that exactly decreases a given Lyapunov/energy function, thereby preserving stability by construction [42]. Related effective discretization schemes based on the Runge-Kutta (RK) family can preserve the Lyapunov decrease while reducing computational cost. In particular, reduced RK (RRK) methods simplify the standard RK update by removing numerically insignificant terms, maintaining the same stability properties with fewer operations [43]. For homogeneous systems, Lyapunov-based discretizations can preserve exponential, FT, and FxT convergence rates and the Lyapunov function across the continuous-to-discrete transition [44]. More generally, a unified framework for translating Lyapunov arguments into discrete-time updates was proposed in [45], ensuring that discrete algorithms inherit the same Lyapunov guarantees as their continuous-time counterparts. FT and FxT guarantees for convex optimization, including proximal and minimax variants, have been established in continuous-time and analyzed under forward-Euler discretization (with small step sizes) [46]. These results suggest that discrete implementations should be guided by Lyapunov inequalities rather than tied to a specific numerical scheme.
@@ Appendix A Proofs
Strong convexity of $J$ implies the objective $J$ is strictly convex; affine equality constraints and convex inequalities define a convex feasible set. Under Slater, KKT conditions are necessary and sufficient for optimality; under LICQ and strong convexity, the primal solution $\mathbf{x}^{\star}$ is unique. Moreover, LICQ ensures uniqueness of the Lagrange multipliers $(\boldsymbol{\lambda}^{\star},\boldsymbol{\mu}^{\star})$ associated with $\mathbf{x}^{\star}$ [31, Thm. 11.12]. Hence, the KKT point is unique.
Let $V(\mathbf{z}):=\tfrac{1}{2}\|\mathbf{S}(\mathbf{z})\|^{2}$. Then $\nabla V(\mathbf{z})=\nabla\mathbf{S}(\mathbf{z})^{\top}\mathbf{S}(\mathbf{z})$. We claim that any critical point of the OLF satisfies $\mathbf{S}(\mathbf{z})=\mathbf{0}$. This is precisely Assumption III.2.
Let $\mathbf{z}=(\mathbf{x},\boldsymbol{\lambda},\boldsymbol{\mu})$ and suppose $\nabla V(\mathbf{z})=\mathbf{0}$. Writing the stationarity of $V$ with respect to $(\boldsymbol{\mu},\boldsymbol{\lambda},\mathbf{x})$ yields:

### v5

@@ V-B Convex-Concave Case
By Sion’s minimax theorem [sion_general_1958], a saddle point $(\mathbf{x}^{\star},\mathbf{y}^{\star})$ exists. Along the dynamics (55), the derivative of $V$ satisfies
@@
Classical minimax dynamics have established asymptotic or exponential convergence under convex-concave assumptions [feijer_stability_2010, cherukuri_saddle-point_2017]. Our framework recovers these results as special cases and further guarantees FT, FxT, or PT convergence by an appropriate choice of the decay law.
@@ V-C Strongly Convex-Strongly Concave Case
Similarly to the constrained optimization case, when strong convexity/concavity holds, we get automatic nonsingularity of Assumption III.2. Moreover, as discussed in Remark IV.7, radial unboundedness of the Lyapunov function is not required, since the saddle-point dynamics are expressed in feedback form and inherently ensure bounded trajectories through the closed-loop structure.
@@
Classical constrained minimax dynamics achieve exponential convergence for convex-concave problems [feijer_stability_2010, cherukuri_saddle-point_2017]. Recent FT/FxT solutions address unconstrained settings [romero_finite-time_2020, garg_fixed-time_2021, ozaslan_exponential_2024]. Theorem V.1 and Corollary V.2 extend FT/FxT/PT guarantees to constrained convex-concave minimax problems.
This section establishes convergence for constrained convex-concave minimax problems according to a chosen convergence rate, extending beyond existing asymptotic or exponential guarantees in the literature. These results illustrate how the OLF framework generalizes across problem classes with shared feasibility constraints. In the next section, we address the GNE seeking problems, where multiple agents interact strategically under such shared constraints.
@@ VI Generalized Nash Equilibrium Seeking Problems
Generalized Nash equilibrium (GNE) seeking problems arise when agents share coupling constraints, such as common resources or safety requirements, while pursuing individual objectives. These problems have been extensively studied from both the variational inequality and optimization perspectives; see the comprehensive survey in [facchinei_generalized_2010]. Classical reformulations based on Nikaido-Isoda functions and their regularized variants [von_heusinger_optimization_2009] provide optimization-based characterizations of (normalized) GNE. Algorithmic approaches have primarily relied on penalty and augmented Lagrangian methods, which embed the constraints into the agents’ costs. Representative contributions include penalty reformulations [facchinei_penalty_2010], augmented Lagrangian frameworks [kanzow_augmented_2016], and, more recently, continuous-time penalty dynamics that establish asymptotic or exponential convergence [sun_continuous-time_2021].
@@ VI-A Problem statement and stationarity vector
Consider $N$ players, where each player $i\in\{1,\dots,N\}$ chooses a decision vector $\mathbf{x}_{i}\in\mathbb{R}^{n_{i}}$ subject to its local constraint set $\mathcal{X}_{i}$. Let $\mathbf{x}_{-i}$ denote the stacked decisions of all other players. Then, each player seeks to minimize its local cost function
$\min_{\mathbf{x}_{i}\in\mathcal{X}_{i}}J_{i}(\mathbf{x}_{i},\mathbf{x}_{-i}),$ (57)
where $J_{i}:\mathbb{R}^{n}\!\to\!\mathbb{R}$ is continuously differentiable and depends on both the player’s own decision and those of the others.
A profile $\bar{\mathbf{x}}=\operatorname{col}(\bar{\mathbf{x}}_{1},\dots,\bar{\mathbf{x}}_{N})$ with $n=\sum_{i}n_{i}$ is a GNE if it is feasible for the shared constraints and no player can reduce its cost by a unilateral deviation that also respects those shared constraints; namely,
@@
with $\mathcal{X}_{i}(\mathbf{x}_{-i}):=\{\mathbf{x}_{i}\mid\mathbf{A}\mathbf{x}=\mathbf{b},\ \mathbf{g}(\mathbf{x})\leq\mathbf{0}\}$ where $\mathbf{A}\in\mathbb{R}^{q\times n}$, $\mathbf{b}\in\mathbb{R}^{q}$, and $\mathbf{g}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ is $C^{1}$. This differs from a standard Nash equilibrium because each player’s feasible set depends on the others via the shared constraints. In convex settings, equilibria for which all players share the same multipliers $(\boldsymbol{\lambda},\boldsymbol{\mu})$ for the joint constraints are called variational GNE (v-GNE); these coincide with solutions of a single KKT system [dreves_solution_2011] and are the natural target for our OLF design.
@@ VI-B Main result
Note that Assumption III.2 holds automatically in this setting. Furthermore, as discussed in Remark IV.7, radial unboundedness of the Lyapunov function is not required, since the primal–dual dynamics are expressed in feedback form and inherently guarantee bounded trajectories and convergence to the v-GNE.
@@
Existence and uniqueness of the variational equilibrium follow from strong monotonicity of $\mathcal{G}$ and convexity of the shared feasible set, as established by variational inequality theory [facchinei_generalized_2010, Sec. 3.2]. By construction of $\mathbf{u}(\mathbf{z},t)$, the OLF $V(\mathbf{z})$ satisfies the chosen decay law, hence $V(\mathbf{z}(t))\!\to\!0$ with the corresponding temporal guarantee, and thus $\mathbf{S}(\mathbf{z}(t))\!\to\!\mathbf{0}$.
@@ VI-C Numerical Example: Cournot GNE with Shared Constraints
We illustrate Thm. VI.2 with a Cournot competition game [martinez-piazuelo_payoff_2022]. Consider $N=4$ firms and $M=2$ markets. Each firm selects a production vector $\mathbf{x}_{k}\in\mathbb{R}^{2}_{\geq 0}$, and the aggregate supply is $C\mathbf{x}$. Market prices follow a linear inverse demand
@@
It is also worth noting that the observation of Remark IV.8 carries over unchanged here: since the Lyapunov function is built directly from the stacked stationarity vector, driving $V(\mathbf{z})$ below a given tolerance automatically certifies that each optimality component (stationarity, feasibility, and complementarity) is satisfied to the same numerical accuracy. This single scalar measure thus provides a unified guarantee for convergence to the v-GNE.
@@ VII Discussion About Discrete-Time Implementation
This section discusses the discrete-time realization of our continuous-time designs. Several recent directions inform the discrete-time design of continuous-time optimization dynamics. Discrete-gradient (DG) methods construct a step that exactly decreases a given Lyapunov/energy function, thereby preserving stability by construction [hernandez-solano_numerical_2023]. Related effective discretization schemes based on the Runge-Kutta (RK) family can preserve the Lyapunov decrease while reducing computational cost. In particular, reduced RK (RRK) methods simplify the standard RK update by removing numerically insignificant terms, maintaining the same stability properties with fewer operations [guedes_preservation_2025]. For homogeneous systems, Lyapunov-based discretizations can preserve exponential, FT, and FxT convergence rates and the Lyapunov function across the continuous-to-discrete transition [sanchez_lyapunov-based_2021]. More generally, a unified framework for translating Lyapunov arguments into discrete-time updates was proposed in [ushiyama_unified_2023], ensuring that discrete algorithms inherit the same Lyapunov guarantees as their continuous-time counterparts. FT and FxT guarantees for convex optimization, including proximal and minimax variants, have been established in continuous-time and analyzed under forward-Euler discretization (with small step sizes) [garg_fixed-time_2023]. These results suggest that discrete implementations should be guided by Lyapunov inequalities rather than tied to a specific numerical scheme.
@@ Appendix A Proofs
Strong convexity of $J$ implies the objective $J$ is strictly convex; affine equality constraints and convex inequalities define a convex feasible set. Under Slater, KKT conditions are necessary and sufficient for optimality; under LICQ and strong convexity, the primal solution $\mathbf{x}^{\star}$ is unique. Moreover, LICQ ensures uniqueness of the Lagrange multipliers $(\boldsymbol{\lambda}^{\star},\boldsymbol{\mu}^{\star})$ associated with $\mathbf{x}^{\star}$ [beck_introduction_2014, Thm. 11.12]. Hence, the KKT point is unique.
Let $V(\mathbf{z}):=\tfrac{1}{2}\|\mathbf{S}(\mathbf{z})\|^{2}$. Then $\nabla V(\mathbf{z})=\nabla\mathbf{S}(\mathbf{z})^{\top}\mathbf{S}(\mathbf{z})$. Furthermore, let $\mathbf{z}=(\mathbf{x},\boldsymbol{\lambda},\boldsymbol{\mu})$ and suppose $\nabla V(\mathbf{z})=\mathbf{0}$. Writing the stationarity of $V$ with respect to $(\boldsymbol{\mu},\boldsymbol{\lambda},\mathbf{x})$ yields: